# MambaMoE: Mixture-of-Spectral-Spatial-Experts State Space Model for Hyperspectral Image Classification

### v1

@@ 3.3. Uncertainty-Guided Corrective Learning
(11) $\mathbf{A}_{k}=\frac{\exp(\mathbf{S}_{k})}{\sum\limits_{j=1}^{n_{class}}\exp(\mathbf{S}_{j})},$
@@
(12) $\mathbf{P}_{r,c}=\max\limits_{k\in\{1,\dots,n_{class}\}}\mathbf{A}_{k,r,c},\quad\mathbf{P}_{r,c}\in[0,1].$
@@
(13) $\mathbf{U}_{r,c}=-\log(\mathbf{P}_{r,c}+\epsilon)\cdot\mathbf{P}_{r,c},$
@@
(14) $M_{r,c}=\begin{cases}1&\text{if }r_{r,c}<\mathbf{U}_{r,c}\\ 0&\text{if }r_{r,c}\geq\mathbf{U}_{r,c},\end{cases}$
@@
3.4. Training Loss
In this study, since image-level HSI classification is formulated as a multi-class segmentation task, we adopt the cross-entropy loss function as the objective criterion. The total training loss comprises two components: (1) the losses computed at each intermediate stage using the sampled supervision labels, and (2) the loss from the final prediction, which is computed against the original ground-truth labels. All loss terms are weighted equally, with a coefficient of 1.
4. Experiments
4.1. Datasets & Experimental Settings
We utilize three hyperspectral benchmark datasets: Pavia University, Houston, and Whu-HanChuan to evaluate the performance of the proposed method. The dataset descriptions and detailed experimental settings are provided in the supplementary material.
Table 1. Performance contribution of different module on three benchmark datasets (reported in OA (%). Best results are highlighted in Bold).
@@ 4.1. Datasets & Experimental Settings
4.2. Ablation Study
Network Blocks. To validate the effectiveness of the proposed MoMEB and UGCL, we conduct ablation studies on three HSI datasets. The baseline model includes only multi-scale feature extraction and stage-wise fusion, without the incorporation of either MoMEB or UARB. We evaluate the impact of integrating MoMEB and UARB individually and jointly. As shown in Table 1, both modules contribute significantly to performance improvement. On the Pavia University dataset, incorporating MoMEB improves the baseline OA by 2.33%, while UARB yields a 1.25% increase. When both modules are applied together, the model achieves the highest OA of 95.20%, representing an overall improvement of 3.67% over the baseline. These findings highlight the individual merits of MoMEB and UARB, as well as their complementary synergy when combined.
Table 2. The performance of spectral-spatial expert mechanism on three benchmark datasets (reported in OA (%)). The best results are highlighted in Bold. “w/o SRE/SSE” represents the absence of the routed spatial expert or shared spectral expert, while “w/ SRE & SSE” indicates that dynamic spectral-spatial experts are employed.
@@ 4.2. Ablation Study
Internal Design of DSSEM. To further investigate the internal architecture of the DSSEM in the proposed MoMEB module, we conduct an ablation study by isolating the contributions of the SRE and SSE. This analysis aims to evaluate the individual and combined impact of these components on classification performance. As shown in Table 2, the removal of either SRE or SSE results in a consistent decline in accuracy across all three HSI datasets, highlighting the importance of jointly leveraging spatial and spectral contextual modeling. For example, on the Pavia University dataset, the full DSSEM achieves an OA of 95.20%. In contrast, employing only SRE or only SSE leads to OAs of 94.24% and 93.17%, corresponding to performance drops of 0.96% and 2.03%, respectively. These findings affirm that the complementary interaction between SRE and SSE is crucial for effective spectral-spatial representation learning, thereby validating the design rationale of the DSSEM.
Number of Selected Experts in SRE. After MoMEB and DSSEM, we further investigate the optimal configuration of the SRE, specifically focusing on the impact of varying the number of top-$k$ selected experts during inference. To this end, experiments were conducted on all three benchmark datasets to evaluate classification performance under different $k$ values. As reported in Table 3, the OA consistently improves as $k$ increases, with the best performance achieved at $k=3$ across all datasets. This performance gain is attributed to the enhanced spatial context modeling facilitated by incorporating a broader set of directional experts, which proves beneficial for handling the complexity and heterogeneity of ground objects in hyperspectral scenes. However, a noticeable decline in OA is observed when $k=4$, suggesting a diminishing return beyond a certain level of expert aggregation. This drop can be explained by the introduction of redundant or less relevant features, which potentially dilute the discriminative capacity of the learned representations. Moreover, when $k=4$, the model effectively reduces to a fully weighted combination of all experts, thereby losing the sparsity and dynamic routing characteristics fundamental to the MoE paradigm. These results underscore that selecting the top-3 experts during inference offers the most effective trade-off between spatial diversity and feature relevance, thereby yielding the highest classification accuracy.
Table 3. Effects of number of top-k experts on the three benchmark datasets (reported in OA (%)). The best results are highlighted in Bold.
@@
Table 4. Quantitative classification results of the three benchmark datasets. Best results are shown in bold.
Type Methods Pavia University Houston Whu-HanChuan
OA $\uparrow$ Kappa $\uparrow$ AA $\uparrow$ OA $\uparrow$ Kappa $\uparrow$ AA $\uparrow$ OA $\uparrow$ Kappa $\uparrow$ AA $\uparrow$
CNN-based SSFCN 73.89±1.00 66.89±1.22 78.37±0.67 79.01±0.96 77.34±1.04 79.46±1.05 75.53±1.42 71.81±1.60 72.72±1.56
SACNet 71.25±1.55 64.38±1.86 78.77±2.13 62.74±1.72 59.87±1.86 66.10±1.79 77.47±2.29 74.05±2.60 76.09±2.70
FcontNet 82.63±3.09 77.68±3.88 83.54±2.74 70.74±4.32 68.47±4.64 74.78±3.66 81.00±6.23 78.04±7.09 78.68±5.71
Transformer-based SpectralFormer 70.53±3.78 63.00±4.27 77.69±1.65 77.85±1.00 76.10±1.07 78.75±1.39 76.34±1.83 72.67±2.00 72.16±1.07
GAHT 90.46±2.66 87.67±3.29 93.24±1.11 89.43±1.16 88.58±1.25 90.83±0.87 87.12±1.69 85.05±1.89 85.94±1.24
CSIL 84.85±2.79 80.4±3.41 86.02±1.78 70.94±1.83 68.66±1.96 74.89±1.64 82.89±1.45 80.29±1.65 86.73±1.22
MASSFormer 91.32±3.96 88.81±4.84 94.45±1.58 89.83±1.59 89.01±1.72 91.30±1.50 90.75±1.98 89.23±2.27 90.22±1.61
DSFormer 92.81±2.20 90.66±2.78 95.38±1.15 88.80±1.72 87.90±1.86 90.32±1.53 91.68±1.28 90.31±1.47 91.22±0.84
Mamba-based SS-Mamba 91.58±2.35 89.13±2.92 95.76±1.27 89.83±2.22 89.01±2.38 91.23±1.68 83.14±4.48 80.44±5.16 81.21±4.47
MambaHSI 92.08±0.70 89.57±0.90 91.11±0.55 86.96±0.52 85.91±0.56 88.70±0.38 91.27±0.25 89.84±0.30 90.89±0.65
HyperMamba 79.89±1.76 74.21±2.05 85.09±0.83 85.04±0.80 83.86±0.87 87.03±0.97 87.58±0.70 85.55±0.79 85.21±0.18
S2Mamba 68.98±5.99 61.57±6.31 79.02±2.23 75.28±2.92 73.33±3.15 79.28±2.22 79.83±2.39 76.66±2.66 75.86±1.79
MambaLG 81.53±1.68 76.41±2.06 87.54±1.38 86.89±2.10 85.83±2.26 87.08±1.71 79.45±1.26 76.23±1.40 77.23±1.30
MambaMoE (Ours) 95.20±1.31 93.74±1.68 97.66±0.55 90.94±0.69 90.21±0.74 92.23±0.62 92.67±1.22 91.42±1.43 91.28±1.70
Figure 4. Visualization of the classification maps on Pavia University dataset. (a) False-color image. (b) Ground truth. (c) SSFCN. (d) SACNet. (e) FcontNet. (f) SpectralFormer. (g) GAHT. (h) CSIL. (i) MASSFormer. (j) DSFormer. (k) SS-Mamba. (l) MambaHSI. (m) HyperMamba. (n) S2Mamba. (o) MambaLG. (p) MambaMoE.
4.3. Performance Comparison and Analysis
In this subsection, we conduct a comprehensive comparison between the proposed MambaMoE and existing state-of-the-art (SOTA) HSI classification methods from both quantitative and qualitative perspectives. The compared methods fall into three main types: (1) CNN-based approaches, including SSFCN (Xu et al., 2020), SACNet (Xu et al., 2021), and FcontNet (Wang et al., 2022); (2) Transformer-based methods, including SpectralFormer (Hong et al., 2022), GAHT (Mei et al., 2022), CSIL (Yang et al., 2023), MASSFormer (Sun et al., 2024), and DSFormer (Xu et al., 2025); and (3) Mamba-based architectures, including SS-Mamba (Huang et al., 2024), MambaHSI (Li et al., 2024b), HyperMamba (Liu et al., 2024b), S2Mamba (Wang et al., 2025), and MambaLG (Pan et al., 2025).
Quantitative Results. The quantitative results are presented in Table 4, highlighting the comparative performance of various state-of-the-art methods. Since MambaMoE integrates image-level feature extraction with dynamic spectral-spatial context modeling through adaptive directional feature extraction and sparse expert activation. Moreover, the UGCL strategy enhances model learning on challenging regions by enforcing stage-wise supervision, thereby mitigating the limitations of Mamba’s inherently sequential modeling architecture. As a result, MambaMoE consistently outperforms all competing methods, including most Mamba-based approaches that rely on patch-level classification, which typically generate representations with limited receptive fields and struggle to capture the holistic semantics of remote sensing scenes. Even compared with MambaHSI, which also utilizes image-level input and achieves an OA of 92.08% on the Pavia University dataset
Qualitative Results We further present the classification maps in Fig. 4 to visually evaluate the performance of various methods. As illustrated, the proposed MambaMoE delivers superior classification results, characterized by smooth and coherent land-cover structures, well-defined boundaries, and minimal salt-and-pepper noise, closely aligning with the ground truth. Notably, the areas highlighted by the black boxes reveal that methods such as SSFCN, SpectralFormer, HyperMamba, and MambaLG suffer from considerable salt-and-pepper artifacts. Additionally, models like SACNet, FcontNet, DSFormer, and MambaHSI exhibit issues such as misclassification and blurred boundaries. In contrast, MambaMoE accurately delineates the contours of the “Bitumen” class and achieves precise classification of categories like “Bricks” and “Asphalt,” preserving both spatial consistency and semantic clarity.
Robustness with fewer training samples. To rigorously evaluate the stability of the proposed MambaMoE architecture, we conducted additional experiments under more challenging conditions using fewer training samples. Specifically, in addition to the standard experimental setup, we created training subsets containing only 5 and 10 samples per class for each of the three benchmark datasets. Subsequently, we compared the OA of MambaMoE with several strong baselines: DSFormer, S2Mamba and MambaHSI, which were selected based on their competitive performance in Table 4, as shown in Fig. 5. The experimental results reveal that MambaMoE consistently outperforms all competitors, even with severely limited supervision. Notably, the drop in performance remains minimal, demonstrating that MambaMoE retains effectiveness despite reduced training data. These findings underscore the robustness, generalization capability, and practicality of MambaMoE in real-world scenarios where labeled hyperspectral data is often scarce.
Figure 5. Comparison of classification performance with fewer training samples: (a) Pavia University dataset. (b) Houston dataset. (c) Whu-HanChuan dataset.
Computation Cost Analysis. We conducted a comprehensive evaluation of the proposed MambaMoE against several Transformer- and Mamba-based baselines, focusing on four critical metrics: model parameters, floating-point operations (FLOPs), inference time, and OA. The comparative results are summarized in Table 5. The findings demonstrate that MambaMoE achieves a highly favorable balance between classification accuracy and computational efficiency. It consistently delivers the highest OA among all compared models while incurring the lowest FLOPs and exhibiting the fastest inference time. Although MambaMoE does not have the absolute smallest number of parameters, the difference is minimal (approximately 0.1M) and thus negligible in practical terms. For instance, on the Pavia University dataset, MambaMoE attains an OA of 95.20%, with only 6.71M FLOPs and an inference time of 0.10 seconds. These performance benefits largely stem from the sparse activation mechanism employed in the MoE architecture, where only a selective subset of expert modules is activated per forward pass, significantly reducing redundant computation. Additionally, the use of image-level inputs enables parallel prediction across a large number of samples, further improving inference efficiency. In summary, MambaMoE delivers state-of-the-art performance with minimal computational overhead, making it especially well-suited for deployment in resource-constrained or real-time hyperspectral remote sensing applications.
Table 5. Comparison of model parameters (M), FLOPs (M), inference time (s) and OA (%) across three benchmark datasets.

### v2

@@ 3.3 Uncertainty-Guided Corrective Learning
$\mathbf{A}_{k}=\frac{\exp(\mathbf{S}_{k})}{\sum\limits_{j=1}^{n_{class}}\exp(\mathbf{S}_{j})},$ (11)
@@
$\mathbf{P}_{r,c}=\max\limits_{k\in\{1,\dots,n_{class}\}}\mathbf{A}_{k,r,c},\quad\mathbf{P}_{r,c}\in[0,1].$ (12)
@@
$\mathbf{U}_{r,c}=-\log(\mathbf{P}_{r,c}+\epsilon)\cdot\mathbf{P}_{r,c},$ (13)
@@
$M_{r,c}=\begin{cases}1&\text{if }r_{r,c}<\mathbf{U}_{r,c}\\ 0&\text{if }r_{r,c}\geq\mathbf{U}_{r,c},\end{cases}$ (14)
@@
3.4 Training Loss
In this study, since image-level HSI classification is formulated as a multi-class segmentation task, we adopt the cross-entropy loss function as the objective criterion. The total training loss comprises two components: (1) the losses computed at each intermediate stage using the sampled supervision labels, and (2) the loss from the final prediction, which is computed against the original ground-truth labels. All loss terms are weighted equally, with a coefficient of 1. Therefore the overall loss function of this work $\mathcal{L}$ is defined as follows:
$\mathcal{L}=\sum_{i=1}^{3}\mathcal{L}_{ce}(\mathbf{Q}_{i},\mathbf{A}_{i})+\mathcal{L}_{ce}(\mathbf{GT},\mathbf{A}_{3}),$ (15)
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function, $\mathbf{Q}_{i}$, $\mathbf{A}_{i}$, and GT correspond to the sampled labels, the predicted probability, and the training labels, respectively.
4 Experiment
4.1 Dataset
We employ three benchmark hyperspectral datasets to evaluate the effectiveness of the proposed method. A detailed description is provided below.
  1. (1)
    Pavia University Dataset. This dataset is captured by the Reflective Optics System Imaging Spectrometer (ROSIS) in 2001, this dataset covers the University of Pavia in northern Italy. It includes 103 spectral bands and a spatial resolution of 610 × 340 pixels, encompassing nine land-cover classes.
  2. (2)
    Houston Dataset. This dataset is acquired in 2012 over the University of Houston and its surrounding urban areas, this dataset consists of 144 spectral bands within the 400–1000 nm wavelength range. It has a spatial resolution of 2.5 meters, an image size of 349 × 1905 pixels, and includes 15 semantic categories [36].
  3. (3)
    Whu-HanChuan Dataset. This dataset is introduced by Zhong et al. [37], this dataset was obtained via UAV-borne imaging, offering rich spatial detail with centimeter-level resolution. It features 274 spectral bands within the 400–1000 nm range and a spatial size of 1217 × 303 pixels.
4.2 Experimental Settings
The experiments in this study were performed using the PyTorch framework, with an NVIDIA GeForce RTX 3090 GPU for computation. For the first two datasets, only 15 samples per class were randomly chosen for the training set, while 30 samples per class were selected for the Whu-HanChuan dataset. The remaining labeled samples were used exclusively as the test set. The training process employed the Adam optimizer with a learning rate of $5\times 10^{-4}$ and was conducted for 200 epochs across all datasets. To quantitatively evaluate classification performance, three widely recognized metrics were employed: overall accuracy (OA), average accuracy (AA), and the kappa coefficient ($\kappa$). Each experiment was repeated 10 times with different random seeds, and the results are reported as the mean and standard deviation.
Table 1: Performance contribution of different module on three benchmark datasets (reported in OA (%). Best results are highlighted in Bold).
@@ 4.2 Experimental Settings
4.3 Ablation Study
(1) Network Blocks. To validate the effectiveness of the proposed MoMEB and UGCL, we conduct ablation studies on three HSI datasets. The baseline model includes only multi-scale feature extraction and stage-wise fusion, without the incorporation of either MoMEB or UARB. We evaluate the impact of integrating MoMEB and UARB individually and jointly. As shown in Table 1, both modules contribute significantly to performance improvement. On the Pavia University dataset, incorporating MoMEB improves the baseline OA by 2.33%, while UARB yields a 1.25% increase. When both modules are applied together, the model achieves the highest OA of 95.20%, representing an overall improvement of 3.67% over the baseline. These findings highlight the individual merits of MoMEB and UARB, as well as their complementary synergy when combined.
Table 2: The performance of spectral-spatial expert mechanism on three benchmark datasets (reported in OA (%)). The best results are highlighted in Bold. “w/o SRE/SSE” represents the absence of the routed spatial expert or shared spectral expert, while “w/ SRE & SSE” indicates that dynamic spectral-spatial experts are employed.
@@ 4.3 Ablation Study
Table 3: Effects of number of top-k experts on the three benchmark datasets (reported in OA (%)). The best results are highlighted in Bold.
@@
(2) Internal Design of DSSEM. To further investigate the internal architecture of the DSSEM in the proposed MoMEB module, we conduct an ablation study by isolating the contributions of the SRE and SSE. This analysis aims to evaluate the individual and combined impact of these components on classification performance. As shown in Table 2, the removal of either SRE or SSE results in a consistent decline in accuracy across all three HSI datasets, highlighting the importance of jointly leveraging spatial and spectral contextual modeling. For example, on the Pavia University dataset, the full DSSEM achieves an OA of 95.20%. In contrast, employing only SRE or only SSE leads to OAs of 94.24% and 93.17%, corresponding to performance drops of 0.96% and 2.03%, respectively. These findings affirm that the complementary interaction between SRE and SSE is crucial for effective spectral-spatial representation learning, thereby validating the design rationale of the DSSEM.
@@
Table 4: Quantitative classification results of the Pavia University dataset. Best results are shown in bold.